OpenStudy (anonymous):

Probability: In a certain road, 1/3 of the houses have no newspapers delivered. If 1/4 have a national paper delivered and 3/5 have a local paper delivered, what is the probability that a house chosen at random has both? Assuming that national paper = n, and local = L, I did P(N n L), but the problem is, since this is a mutually exclusive event, won't the P(N n L) = 0?

OpenStudy (anonymous):

I confused whether I should use this formula: P(A u B) = P(A) + P(B) - P(A n B), or this: P(A u B) = P(A) + P(B). how do I even get P(A n B)???

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):



OpenStudy (tkhunny):

1 - 1/3 = 2/3 2/5 + 1/4 - "The Little Piece" = 2/3

OpenStudy (anonymous):

3/5 how 2/5?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Show a little fortitude and just fix errors when you encounter them. I can take it. 3/5 + 1/4 - "The Little Piece" = 2/3 Solve for the little piece.

OpenStudy (anonymous):

p(a union b)=2/3

OpenStudy (anonymous):

91/60?

OpenStudy (tkhunny):

Let's see. 91/60 > 1 -- That doesn't seem right. Doesn't it have to be less than 1/4? 3/5 + 1/4 - "The Little Piece" = 2/3 12/20 + 5/20 - "The Little Piece" = 2/3 17/20 - "The Little Piece" = 2/3 17/20 - 2/3 = "The Little Piece" 51/20 - 40/3 = "The Little Piece" Ah! I see the problem. Can you fix it?

OpenStudy (anonymous):

3/5+1/4+2/3=is it right?

OpenStudy (tkhunny):

Did you look at what I showed you? You have a sign wrong. The answer must be less than 1/4. If you get something greater than 1/4, something is very wrong.

OpenStudy (anonymous):

11/60

OpenStudy (anonymous):

it shows in your equation above 3/5 + 1/4 - "The Little Piece" = 2/3

OpenStudy (anonymous):

if 2/3 from right hand side comes lefthand side you will get + sign

OpenStudy (anonymous):

51/20 - 40/3 = "The Little Piece"

OpenStudy (anonymous):

51-40/60

OpenStudy (anonymous):

sorry to budge in, but I keep getting 11/60 for answer?

OpenStudy (tkhunny):

Typo Again! 51/60 - 40/60 = "The Little Piece" Notation is wrong. 51-40/60 = 51 - (40/60) and is NOT what you intend. Okay, is 11/60 correct?

OpenStudy (anonymous):

no. its 11/20 D:

OpenStudy (tkhunny):

Let's see: 1/3 + 3/5 + 1/4 - 11/20 = 19/30 Nope. That is NOT Unity (1) Let's see: 1/3 + 3/5 + 1/4 - 11/60 = 60/60 = 1 Yup. That is Unity (1)